The height of a triangle sail is 5 feet more than twice it's base. If the sail is made of 125 sq ft of fabric what is height and base?
Answer by TimothyLamb(4379)

h = 2b + 5
a = (1/2)bh
---
125 = (1/2)bh
125 = (1/2)b(2b + 5)
250 = 2bb + 5b
2bb + 5b - 250 = 0
---
2x^2 + 5x - 250 = 0
---
the above quadratic equation is in standard form, with a=2, b=5, and c=-250
---
to solve the quadratic equation, plug this:
2 5 -250
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
the roots of the quadratic are:
b = 10
b = -12.5
---
the negative root doesn't make sense for a length, so use the positive root:
---
Answer:
b = 10 ft
h = 25 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php